### v1

# Binary Stars Approaching Supermassive Black Holes: Hydrodynamics of Stellar Collisions, Mass Fallback and Partial TDEs

Abstract
When binaries are injected into low-angular-momentum orbits around a central supermassive black hole (SMBH), various outcomes can occur, including binary tidal breakup, double stellar disruptions and stellar collision. We use hydrodynamical simulations to study stellar collisions triggered by binary-SMBH encounters, examining both head-on and grazing collisions in deep ($\beta_{b}=5$) and gentle ($\beta_{b}=0.6$) encounters, where $\beta_{b}$ is the ratio of the binary tidal disruption radius to the binary pericenter distance to the SMBH. Head-on collisions consistently result in appreciable mass loss ($\sim 5\%$) and a single merger remnant. Grazing collisions in deep encounters typically leave two strongly disturbed stars with comparable mass loss, while in gentle encounters, multiple collisions eventually produce a single remnant with minimal mass loss ($\lesssim 1\%$). All merger remnants feature extended envelopes, making them susceptible to partial tidal disruptions when they return to the SMBH. The morphology and orbital energy distribution of collision-induced debris differ significantly from those of tidal disruption event (TDE) debris of single stars. Approximately half of the collision-generated debris falls back onto the SMBH, exhibiting a distinct time evolution of the fallback rate. We suggest that such mass loss and fallback can generate electromagnetic flares that mimic weak TDEs.
Black holes; Binary stars; Tidal disruption; Stellar dynamics
software: PHANTOM (Price et al., 2018), Sarracen (Harris & Tricco, 2023), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011).
1 Introduction
Almost all galaxies harbor a supermassive black hole (SMBH) at their center (see Kormendy & Ho, 2013, for a review), and the surrounding nuclear clusters are expected to have high stellar densities. This unique environment leads to complex dynamical processes and diverse phenomena around the SMBH (see Alexander, 2017, for a review). A large fraction of stars in nuclear clusters are members of stellar binaries. The interactions (i.e., gravitational scatterings) between binaries and single stars tend to destroy relatively wide (or soft) binaries, leaving behind only tightly bound (or hard) binaries (e.g., Heggie, 1975; Hills, 1975; Heggie, 2003). These tight binaries can be scattered into low-angular-momentum orbits toward the SMBH by interactions with other stars. When this happens, the SMBH can tidally breakup the binary, resulting in one star becoming bound to the SMBH and the other being ejected as a hyper-velocity star (HVS) that can escape the galaxy (e.g., Hills, 1988; Yu & Tremaine, 2003; Gould & Quillen, 2003). This process (“Hills mechanism”) has been extensively studied as a significant source of HVSs and tightly bound stars in the Galaxy (e.g., Gualandris et al., 2005; Bromley et al., 2006; Ginsburg & Loeb, 2006; Sesana et al., 2007; Sari et al., 2009; Brown et al., 2018).
When individual stars approach the SMBH on sufficiently tight orbits, they may be disrupted by tidal forces, leading to tidal disruption events (TDEs) (e.g., Rees, 1988; Stone et al., 2020). In recent years, stars bound to the SMBH following binary breakup have been linked to repeating partial TDEs (e.g., Cufari et al., 2022; Somalwar et al., 2023). Such stars have also been associated with quasi-periodic eruptions (QPEs), which are periodic bursts of accretion activities that are thought to result from the interaction between bound stars and the SMBH or the accretion disk around the SMBH (e.g., Lu & Quataert, 2023; Linial & Sari, 2023; Zhou et al., 2024, 2024; Yao et al., 2025).
Taking the finite size of stars into account reveals additional phenomena beyond simple binary breakups. For example, double stellar tidal disruptions can occur when both components of a binary are torn apart by the SMBH’s tidal forces (e.g., Mandel & Levin, 2015; Bradnick et al., 2017; Bonnerot & Rossi, 2019). Even when the stars are not tidally disrupted, stellar collisions may happen during binary-SMBH encounters when the stars are brought close to each other due to the gravitational influence of the SMBH (e.g., Ginsburg & Loeb, 2007; Antonini et al., 2010, 2011; Bradnick et al., 2017; Yu & Lai, 2024). Such a collision results in stellar merger and mass loss, which may lead to accretion flares onto the SMBH. The stellar merger can also lead to the formation of exotic stars in nuclear star clusters.
In Yu & Lai (2024) (hereafter Paper I), we used three-body scattering experiments to characterize the key dynamics of binary tidal breakups, double stellar disruptions, and stellar collisions induced by binary-SMBH encounters. We analyzed the collision probabilities of binaries with different orbital properties and concluded that the collision probability can reach a few to 10’s percent as the binary binary’s pericenter distance (relative to the SMBH) varies from outside the tidal sphere to within it. Additionally, we found that the impact velocities of the colliding stars are comparable to the escape velocity of the star.
In this paper, we extend the work of Paper I by performing detailed hydrodynamical simulations of stellar collisions generated by binary-SMBH encounters. Our goal is to address several questions that cannot be resolved by point-mass scattering experiments alone: What are the outcomes of such collisions under various conditions - do they result in the formation of a single merger remnant, or do they simply leave two stars highly perturbed? What is the post-collision structure of the stars? How much mass is lost during the collision, and could the fallback of this mass onto the SMBH produce a distinctive observable signature?
Our paper is organized as follows. In Section 2, we describe the setup of our numerical simulations, including the initial conditions and methodology. In Section 3, we present the results from our representative simulations. Finally, in Section 4, we summarize our findings and comment on the results of Paper I based on our hydrodynamical simulations.
2 Method and Setup
We performed simulations of stellar collisions during binary-SMBH encounters using the smoothed particle hydrodynamics (SPH) code Phantom (Price et al., 2018).
Following Paper I, we adopt three key dimensionless ratios to characterize the encounters, defined as follows:
  1. 1.
    $\beta_{b}\equiv r_{\rm tide}^{b}/r_{p}$, where $r_{p}$ is the pericenter distance of the binary relative to the SMBH, and
    $r_{\rm tide}^{b}\equiv a_{b}\Bigl{(}\frac{M_{\rm BH}}{m_{12}}\Bigr{)}^{1/3}$ (1)
    is the tidal radius of the binary. Here $a_{b}$ is the binary’s semi-major axis, $M_{\rm BH}$ is the SMBH mass, and $m_{12}$ is the total mass of the binary system.
  2. 2.
    $\beta_{\star}\equiv r_{\rm tide}^{\star}/r_{p}$, where $r_{\rm tide}^{\star}$ is the tidal radius of an individual star in the binary, defined as
    $r_{\rm tide}^{\star}\equiv R_{\star}\Bigl{(}\frac{M_{\rm BH}}{m_{\star}}\Bigr{)}^{1/3}.$ (2)
    Here $R_{\star}$ and $m_{\star}$ are the radius and mass of the star, respectively.
  3. 3.
    $\alpha\equiv a_{b}/R_{\rm col}$, where $R_{\rm col}$ is the separation between the two stars at the moment of collision. For an equal-mass binary ($m_{1}=m_{2}$), $R_{\rm col}=2R_{\star}$.
2.1 Simulation Setup
Throughout this paper, we consider circular equal-mass binary, where each star modeled as a $\gamma=5/3$ polytrope, with mass $m_{\star}=1M_{\odot}$ and radius $R_{\star}=1R_{\odot}$. The binaries in our simulations are then created by shifting the centers of two stellar models; each model is relaxed so that it reaches hydrostatic equilibrium. In each simulation, the binary is placed on a Newtonian parabolic orbit around a $10^{6}M_{\odot}$ SMBH from a distance of $(2-3)\,r_{\rm tide}^{b}$. Note that the probability of stellar collision depends on the orbital phase of the binary. To ensure collision, we adjust the initial binary orbital phase by evolving the binary for certain duration before releasing it on the orbit toward the SMBH.
For the penetrating factor $\beta_{b}$, we select $\beta_{b}=0.6$ to represent a gentle encounter ($\beta_{b}<1$) and $\beta_{b}=5$ for a deep encounter ($\beta_{b}>1$). Throughout the paper, we consider initially coplanar, circular binaries with $\alpha=8$ (see Paper I for the effect of binary eccentricity on the collision probability).
We have performed simulations with $2\times 10^{5}$, $10^{6}$, and $2\times 10^{6}$ SPH particles, and found very little deviations in the simulation results (see Figures 5,11). Unless otherwise noted, all results presented in this paper are from the $10^{6}$ particle runs.
We run each simulation for a few hundred stellar dynamical time ($t_{\rm dyn}\equiv\sqrt{R_{\star}^{3}/Gm_{\star}}$) after the collision to ensure that the energies of the collision remnants have converged and the final collision outcomes are clearly determined.
2.2 Identify Post-collision Outcomes
In Paper I, we adopted the simplest assumption for the collision process, treating it as a perfectly inelastic collision when the separation between the two stars becomes smaller than $R_{\rm col}$, resulting in a single collision remnant. However, in our SPH simulations, we observe that collisions can lead to some degree of mass loss, and it is also possible that the two stellar remnants do not merge. From the wide range of simulations we have carried out, we select four that are representative for detailed presentation in Section 3 (see Table 1).
  1. 1.
    Head-on collision in a deep encounter ($\beta_{b}=5$). The stars collide then the binary reaches the pericenter to the SMBH. If the collision occurs with low relative velocity and a small impact parameter, the stars may merge, forming a new, more massive star.
  2. 2.
    Grazing collision in a deep encounter ($\beta_{b}=5$). In a deep encounter, the binary is disrupted after the pericenter passage. If the stars collide with a large impact parameter without merging, no further collisions occur. In this case, one star remains bound to the SMBH, while the other is ejected, similar to the Hills mechanism. However, unlike a typical binary breakup, both stars may experience violent disturbances during the process.
  3. 3.
    Head-on collision in a gentle encounter ($\beta_{b}=0.6$). In a gentle encounter, the binary survives the pericenter passage. However, the binary can gain a sufficiently large eccentricity, which leads to stellar collisions after the pericenter passage. If the two stars collide with low relative velocity and small impact parameter (a head-on collision), they may coalesce to form a new, more massive star.
  4. 4.
    Grazing collision in a gentle encounter ($\beta_{b}=0.6$). In a grazing collision, where the impact parameter is relatively large, the stars may not merge but instead experience significant perturbations, such as mass exchange or mass loss. In a gentle encounter, the binary survives the pericenter passage, allowing for repeated collisions. This increases the likelihood of a merger, even if the first collision does not result in an immediate merger.
A qualitatively description of these simulation outcomes (e.g., whether a collision occurs or how many remnants exist) is relatively straightforward. However, when performing a quantitative analysis (e.g., to determine the mass loss), careful consideration is required to distinguish SPH particles belonging to the remnant and those contributing to mass loss, particularly due to the exchange of particles between the stars during the collision.
In the case where there is only a single remnant, this analysis is relatively straightforward. We can identify a set of the SPH particles, denoted as $S_{r}$, that can be treated as a coherent body (‘remnant’). The mass, position and velocity of the remnant are
${\boldsymbol{x}}_{\rm CM}=\frac{\sum m_{i}{\boldsymbol{x}_{i}}}{\sum m_{i}},$ (3)
${\boldsymbol{v}}_{\rm CM}=\frac{\sum m_{i}{\boldsymbol{v}_{i}}}{\sum m_{i}},$
$m_{\rm rem}^{(r)}=\sum_{i\in S_{r}}m_{i},$
where $m_{i}$, ${\boldsymbol{x}_{i}}$ and ${\boldsymbol{v}_{i}}$ are the mass, position, and velocity of SPH particle $i$. In our simulations, all particles have equal mass. To simplify the calculation, we sum over all particles when computing the center of mass position and velocity. This is justified because in all our simulations, the mass loss is relatively small ($\lesssim 5\%$). Moreover, the mass lost particles are nearly symmetric relative to the remnant’s center of mass, thus having a negligible impact on the results.
Our method for identifying the post-encounter remnant is based on the Bernoulli constant. For each SPH particle $i$ belonging to the remnant, a specific enthalpy is calculated according to:
$h_{i}^{(r)}=\frac{1}{2}({\boldsymbol{v}}_{i}-{\boldsymbol{v}}_{\rm CM})^{2}+u_{i}+\frac{P_{i}}{\rho_{i}}+\phi_{i}^{(r)},$ (4)
where $u_{i}$ and $\rho_{i}$ are its specific internal energy and density, $P_{i}$ is the pressure, and $\phi_{i}^{(r)}$ is the gravitational potential excluding the influence of SMBH. Along a streamline with no dissipation, $h_{i}^{(r)}$ should be a constant according to the Bernoulli theorem. Particles with $h_{i}^{(r)}>0$ do not belong to $S_{r}$ they are considered as the mass loss, unbound from the remnant.
If there are two merger remnants, we can use similar algorithm based on the Bernoulli constant (see Li et al., 2021). But here, for simplicity, we adopt a more straightforward approach. First, we determine the center of mass of all SPH particles, which corresponds to the binary’s center of mass. Then, within a certain region around this center (determined by the approximate distance between the two remnants from our analysis), we compute a 2D grid-based column density, ensuring the resolution is finer than $0.1R_{\star}$. Since the orbits in our simulations are all coplanar, we assume that the $z$-coordinate of the center of mass is always zero-a simplification that has been verified to be valid. We then identify the 2D coordinates of the two density peaks as the positions of the two remnants’ centers of mass. This method is efficient; however, its drawback is that we cannot precisely determine the mass loss fraction. Nevertheless, by examining the density profile around the center of mass, we can still obtain an estimate of the mass loss.
2.3 Calculation of Mass Fallback Rate
Once we have determined which SPH particles are mass loss, we can calculate the time it takes for them to fall back to SMBH (to the pericenter) to determine the mass fallback rate $\dot{M}$.
An approximate method of calculating $\dot{M}$ is as follows. Our hydrodynamical simulations allow us to calculate the binding energy $\varepsilon$ of the mass-loss particles relative to the SMBH and to determine their energy distribution $dM/d\varepsilon$. This distribution can be used to determine the mass fallback rate $\dot{M}$ as a function of time through Kepler’s third law:
$\dot{M}(t)=\frac{dM}{d\varepsilon}\frac{d\varepsilon}{dt}=\frac{2\pi}{3}(GM_{\rm BH})^{2/3}\frac{dM}{d\varepsilon}t^{-5/3}.$ (5)
This approach is highly effective for calculating the mass fallback rate in TDEs, as the star is disrupted at pericenter, resulting in a narrow angular momentum distribution. Consequently, nearly all bound particles must complete a full orbit before returning to pericenter.
However, the above approach for calculating $\dot{M}$ does not always work well for stellar collisions considered here. In deep encounters, collisions typically occur near the pericenter, but in gentle encounters, collisions tend to happen farther away, making Eq. (5) less accurate. Additionally, the mass loss caused by stellar collisions can have a broad angular momentum distribution, implying that the particles do not necessarily need to complete an entire orbit to return to the pericenter. Thus, it is important to analyze the trajectory of each particle individually to accurately determine the mass fallback rate $\dot{M}$.
To compute the pericenter distance $r_{p}$ and fallback time $t_{\text{fb}}$ for each SPH particle, we analyze the orbital parameters derived from the position $\mathbf{r}$ and velocity $\mathbf{v}$ of the particle. The specific angular momentum vector is $\mathbf{h}=\mathbf{r}\times\mathbf{v}$ and the specific orbital energy (relative to the SMBH) is
$\varepsilon=\frac{v^{2}}{2}-\frac{GM_{\rm BH}}{r}.$ (6)
The orbital eccentricity, semi-major axis and semi-latus rectum are then given by
$e=\sqrt{1+\frac{2\varepsilon h^{2}}{(GM_{\rm BH})^{2}}},\quad a=-\frac{GM_{\rm BH}}{2\varepsilon},\quad p=\frac{h^{2}}{GM_{\rm BH}}.$ (7)
The pericenter distance is then
$r_{p}=\frac{p}{1+e}.$ (8)
For bound orbits ($e<1$), the eccentric anomaly ${\cal{E}}$ and the mean anomaly ${\cal{M}}$ are
$\cos{\cal{E}}=\frac{e+\cos\theta}{1+e\cos\theta},\quad{\cal{M}}={\cal{E}}-e\sin{\cal{E}},$ (9)
and the fallback time is
$t_{\text{fb}}=\sqrt{\frac{a^{3}}{GM_{\rm BH}}}(2\pi-{\cal{M}}).$ (10)
For hyperbolic orbits ($e>1$), the hyperbolic anomaly ${\cal{H}}$ and the mean anomaly ${\cal{M}}$ are
$\cosh{\cal{H}}=\frac{e+\cos\theta}{1+e\cos\theta},\quad{\cal{M}}=e\sinh{\cal{H}}-{\cal{H}},$ (11)
with the fallback time given by (if it can reach pericenter)
$t_{\text{fb}}=\sqrt{\frac{(-a)^{3}}{GM_{\rm BH}}}{\cal{M}}.$ (12)
In practice, we measure $\mathbf{r}$ and $\mathbf{v}$ of the unbound (relative to the remnant) SPH particle at a time after stellar collision when the mass loss fraction has stabilized. We then calculate the fallback times for all mass-loss particles that will return to the SMBH. Finally, we manually add the approximate time from the collision to the calculated fallback time and sum over the mass-loss particles to obtain the final mass fallback rate.
3 Results
Runs $\beta_{b}$ $N$ $r_{12}(R_{\odot})$ $|v_{12}|(\sqrt{GM_{\odot}/R_{\odot}})$ ${b_{12}/r_{12}}$ Number of remnant(s) ${\Delta M/M_{\rm tot}}$
1a 5 $5\times 10^{5}$ $3.749$ $1.155$ $0.0104$ $1$ $0.0495$
1b 5 $1\times 10^{5}$ $3.810$ $1.171$ $0.0625$ $1$ $0.0538$
1c 5 $1\times 10^{6}$ $3.723$ $1.147$ $0.0484$ $1$ $0.0492$
2 5 $5\times 10^{5}$ $3.612$ $1.354$ $0.4885$ $2$ $\sim 0.04$
3 0.6 $5\times 10^{5}$ $3.329$ $1.046$ $0.0285$ $1$ $0.0476$
4 0.6 $5\times 10^{5}$ $2.388$ $1.200$ $0.9085$ $1$ $0.0071$
Table 1: Summary of the SPH simulations. Here $\beta_{b}=r_{\rm tide}^{b}/r_{p}$ (with $r_{\rm tide}^{b}$ the binary tidal radius given by Eq. (1) and $r_{p}$ the pericenter distance), and $N$ refers to the number of SPH particles used to represent each star, $r_{12}$ and $v_{12}$ are the relative separation and velocity just before collision, and $b_{12}\equiv|\vec{v}_{12}\times\vec{r}_{12}|/|\vec{r}_{12}|$ is the impact parameter. We also list the number of remnant(s) and the mass loss fraction $\Delta M/M_{\rm tot}$ (where $M_{\rm tot}=2m_{\star}$) at the end of the simulation. The mass loss fractions are all calculated using the Bernoulli method (see Section 2.2), except for Run 2, where it is estimated from the upper row of Figure 9.
Note that it is not possible to determine the exact moment of collision. In our analysis, we use the first numerical output after the collision occurs as the approximate collision time ($t=0$), and $r_{12}$ and $v_{12}$ are measured at $t\approx-2t_{\rm dyn}$ (with $t_{\rm dyn}\equiv\sqrt{R_{\star}^{3}/Gm_{\star}}$).
Our representative simulations are summarized in Table 1, corresponding to the four cases discussed in Section 2.2. For Run 1, we record the results with different resolutions (i.e., the number of SPH particles) to test the consistency of the results. We also record the relative distance $|\vec{r}_{12}|$ and velocity $|\vec{v}_{12}|$ between the two stars approximately $2t_{\rm dyn}$ (where $t_{\rm dyn}\equiv\sqrt{R_{\star}^{3}/Gm_{\star}}$) before the collision and compute the impact parameter, defined as $b_{12}\equiv|\vec{v}_{12}\times\vec{r}_{12}|/|\vec{r}_{12}|$, to get a rough idea of whether the collision is head-on ($b_{12}/r_{12}\ll 1$) or grazing ($b_{12}/r_{12}\sim 1$). Additionally, at the end of the simulation, we record the number of merger remnants and the mass loss fraction computed using the method described in Section 2.2. Note that when two merger remnants are present, we use the mass radial profile centered on each remnant’s centroid to estimate the amount of mass loss.
Figure 1: The density distribution plots of Run 1a at different times for the $z=0$ slice (the orbital plane). The density values in the plot are in visualization code units. Note that the maximum value of the color bar does not represent the actual maximum value in the plot, but is instead chosen for better visualization of the debris stream’s overall structure. In the bottom row, the small inset square shows a zoomed-in version of the region inside the dashed box from the larger plot, with the color bar adjusted for better display of the central merger remnant. The white “$\times$” marks the position of the SMBH. Here, $t=0$ approximately corresponds to the moment of collision.
Figure 2: The same as Figure 1, but for Run 2.
Figure 3: The same as Figure 1, but for Run 3.
Figure 4: The same as Figure 3, but for Run 4.
Figures 1-4 show the density distribution plot at the $z=0$ slice (the orbital plane) for Runs 1-4. These simulations show that, for both head-on collisions (Run 1 & 3) and the grazing collision in gentle encounter (Run 4), the stellar collisions produce a single merger remnant. In contrast, for the grazing collision in the deep encounter (Run 2), the collision leads to violent perturbations of the two stars but does not result in a merger, leaving behind two remnants instead.
In addition to the remnants, stellar collisions also generate a significant amount of mass-loss debris, which can fall back onto the SMBH. Notably, the resulting “debris clouds” from these events are distinct from the slender TDE streams typically seen in TDE simulations (e.g., Evans & Kochanek, 1989; Coughlin et al., 2016; Stone et al., 2020).
We now examine in detail the properties of the remnants for the four representative simulation runs, along with the orbital characteristics of the mass-loss debris and the fallback rate onto the central SMBH.
3.1 Merger Remnant(s)
Figure 5: Mass loss fraction as a function of time for Runs 1 & 3. The lines with the $\bullet$” markers of different colors represent results with different particle numbers ($N=5,1,10\times 10^{5}$ per star, Runs 1a, 1b, 1c). The line with “$\times$” markers represents the result for Run 3. Note that $t=0$ roughly indicates the moment of collision. $M_{\rm tot}$ represents the total mass of the initial binary.
In Section 2.2, we describe how we identify the SPH particles that belong to the merger remnant(s). Figure 5 presents the results of precisely calculating the mass loss fraction for two head-on collisions using the Bernoulli method. It is evident that the fates of most particles - whether they remain part of the merger remnant or are lost - are determined quickly after the collision. This occurs more rapidly in collisions during gentle encounters compared to those in deep encounters, as the higher impact velocity in deep encounters results in a more violent interaction among fluid elements, leading to a slower mass loss process.
We have conducted a convergence test for head-on collisions in deep encounters using simulations with varying particle numbers (Runs 1a, 1b, 1c). The results demonstrate that as the particle number increases, the mass loss fraction quickly converges.
For head-on collisions in both deep and gentle encounters, the mass loss fraction stabilizes at approximately 5%.
Figure 6: The density profile of the merger remnant (left column) and the fraction of the total mass enclosed within a given radius relative to the total mass of the merger remnant (right column) are shown for Run 1a (upper row) and Run 3 (lower row). The solid lines represent the profiles at different times after the collision, while the dashed line corresponds to a $2M_{\odot}$ $\gamma=5/3$ polytrope with radius determined by the mass-radius relation $R_{\star}\propto M_{\star}^{0.8}$, which also approximates the structure of a main-sequence star. The insets in the right panels highlight the mass distribution at the outskirts of the remnant.
In addition to the mass loss, we are also interested in the structural properties of the merger remnant. Clearly, not all particles settle down immediately after the collision, and some asymmetry exists within the merger remnant during the short post-collision phase. For now, we ignore the effects of this asymmetry. Starting from the remnant’s center of mass, we calculate the density $\rho_{\rm rem}$ at different radius $r_{\rm rem}$ by dividing the mass and volume of spherical shells. We also compute the fraction of the total mass enclosed within a given radius $M_{\rm rem}(<r_{\rm rem})$ relative to the total mass of the merger remnant, as determined using the Bernoulli method. Figure 6 illustrates the structure of the merger remnant following the head-on collision (Runs 1 & 3).
It can be seen that the main part of the remnant (accounting for 90% of the total mass) stabilizes within a few $t_{\rm dyn}$, while the remaining 10% of the mass undergoes small change over hundreds of $t_{\rm dyn}$. The outermost layers of the envelope continue to expand, and in the case of deep encounter (Run 1), the density in the region $r_{\rm rem}\gtrsim 5R_{\odot}$ shows a noticeable increase over time, and 99% of the total mass is confined within a radius of $15R_{\odot}$, which continues to grow slowly. In contrast, for the gentle encounter (Run 3), 99% of the mass is contained within $12R_{\odot}$ and remains very stable.
Compared to a typical $2M_{\odot}$ main-sequence star (modeled here as a $\gamma=5/3$ polytrope, with its radius determined by the mass-radius relation $R_{\star}\propto M_{\star}^{0.8}$; represented by the dashed line in Figure 6), the merger remnant exhibits a more concentrated core and an extremely extended envelope. Over 5% of the total mass is distributed in the envelope at $r_{\rm rem}>5R_{\odot}$.
We can also calculate the total energy of the remnant based on the SPH simulation output. Taking the head-on collision in the deep encounter (Run 1a) as an example, we compute the kinetic energy $E_{k,\rm rem}$ relative to the center of mass, gravitational potential energy $E_{p,\rm rem}$, and internal energy $U_{\rm rem}$ for all SPH particles belonging to the remnant. These components were then used to determine the total energy $E_{tot,\rm rem}$ of the remnant. The results are presented in Figure 7.
Figure 7: Energy of the merger remnant in Run 1a as a function of time. The different solid lines represent the kinetic energy $E_{k,\rm rem}$ (relative to the remnant center of mass), gravitational potential energy $E_{p,\rm rem}$, internal energy $U_{\rm rem}$, and total energy $E_{\rm tot,\rm rem}$.
Consistent with the mass loss fraction, the various energy components of the star quickly reach a steady state, and the total energy of the remnant remains negative throughout.
For the grazing collisions (Runs 2 & 4), the larger impact parameters typically prevent the two stars from merging during the first collision. In the case of the deep encounter (Run 2), the tidal forces from the SMBH have already disrupted the binary energy-wise. After the first collision, the two stars do not collide again but instead undergo a process resembling the standard binary tidal breakup: one star becomes bound to the SMBH’s orbit, while the other is ejected. However, both stars have undergone significant perturbations, leaving their structures vastly different from those of typical main-sequence stars.
Figure 8: Separation between the two stars in Run 4 as a function of time, where $t=0$ marks the starting point of the simulation. The orange dashed line indicates the Hill radius $r_{\rm Hill}$ of the star.
In contrast, in the gentle encounter (Run 4), Figure 8 shows that the binary separation remains smaller than the Hill radius $r_{\rm Hill}$ of the star for most of the time, where $r_{\rm Hill}\equiv r_{1}(M_{\star}/M_{\rm BH})^{1/3}$ (with $r_{1}$ the distance of the star to the SMBH). The two stars remain gravitationally bound to each other throughout the process. Consequently, although the stars do not merge immediately after the first collision, significant dissipation of their relative orbital energy during the collision makes subsequent collisions inevitable. These repeated collisions eventually lead to the formation of a single merger remnant. In our Run 4, the binary undergoes three collisions before finally merging into a single remnant.
Figure 9: Same as Figure 6, but for Runs 2 and 4. Note that the “bump” in the blue line in the upper left panel represents the other remnant. When two remnants are present, the exact remnant mass cannot be easily determined. Therefore, we calculate the fraction of the total mass enclosed within a given radius relative to the initial total mass of the system (which is $1M_{\odot}$ for Run 2 and $2M_{\odot}$ for Run 4).
We also analyze the property of the merger remnant following a grazing collision (Runs 2 & 4), with the results shown in Figure 9. Unlike the cases of head-on collisions, for the deep encounter, we select the remnant that is ultimately bound to the SMBH and compared it to a $1M_{\odot}$ $\gamma=5/3$ polytrope. Additionally, when calculating the fraction of mass enclosed within different radii, we no longer use the merger remnant’s total mass determined via the Bernoulli method. Instead, we directly adopt the initial mass ($2M_{\odot}$ for the gentle encounter and $1M_{\odot}$ for the deep encounter). We also include a red dashed line indicating the estimated total mass of the remnant, inferred from the total mass fraction. This suggests that the mass loss fraction is approximately 4% in Run 2 and around 1% in Run 4. The latter is roughly in line with the value obtained using Bernoulli method (see Table 1).
Similar to the results of head-on collisions (Runs 1 & 3), the structure of the merger remnant after a grazing collision is also significantly different from that of the original star. In the deep encounter (Run 2), the remnant resulting from the collision becomes much puffier, with the central density significantly reduced compared to the original star, and a large envelope develops. Ignoring the second peak at $5.5R_{\odot}$ in the upper left panel of Figure 9 (which represents the other remnant), we can observe that more than 90% of the remnant’s mass stabilizes quickly. However, as seen in the upper right panel, the remnant continues to expand slowly.
For the merger remnant formed from multiple collisions in the gentle encounter (Run 4), the stellar structure is similar to that of the head-on collision case: a higher central density with a larger, extended envelope. However, the stabilization of the merger remnant’s structure takes a longer time in this case. Even after several hundred $t_{\rm dyn}$, the density distribution still shows changes, and the star’s radius continues to increase. This is likely due to the higher angular momentum of the merger remnant leading to greater oblateness.
In Paper I, which assumes perfect inelastic collision when the two stars come into contact, we found that most merger remnants formed in deep encounters, as well as a small fraction of those formed in gentle encounters, are bound to the SMBH if the binary originally follows a parabolic orbit relative to the SMBH. We also suggested that one advantage of the merger remnant scenario is that, due to dissipation during the merger process, the remnant stars likely possess very extended envelopes. When these remnants approach the SMBH in subsequent orbits, their extended envelopes are highly susceptible to tidal disruption, leading to repeating partial TDEs that can occur multiple times.
Here, we confirm that the merger remnant indeed has a very extended envelope. Additionally, our SPH simulations with different impact parameters show that not all collisions result in a single merger remnant (we will discuss how this modifies our results from Paper I in Section 4.2). Regarding repeating partial TDEs, our simulations indicate that, regardless of whether there is only one merger remnant or two remnants, the collision process makes partial tidal disruption more likely. In particular, stars bound to the SMBH after binary breakup will have extended envelopes if they experience grazing collisions, making them more prone to partial disruption.
3.2 Mass Loss Debris and Fallback
In Paper I, we suggested the mass loss from stellar collisions induced by binary-SMBH encounters is mass loss will accrete onto the SMBH, generating accretion luminosity or flares even without a stellar TDE. As shown in Figures 1-5, it is indeed the case that, since the collision velocity is typically comparable to the stellar surface escape velocity (see Figure 9 of Paper I), collisions generally result in a mass loss of approximately 5%. This is consistent with the results of stellar collisions in the absence of an SMBH (e.g., Benz & Hills, 1987; Lai et al., 1993; Freitag & Benz, 2005). A key difference between this mass-loss debris and the typical TDE debris is that the former is not produced by tidal forces, and thus does not remain in the same plane or have the very small launch angles characteristic of the TDE debris (which results in an extremely narrow stream). Instead, the mass-loss debris generated by stellar collisions typically has much larger launch angles, and therefore, in terms of morphology, these debris are more like a “cloud” than a “stream”. Similar to TDEs, because our binary is initially on a parabolic orbit relative to the SMBH, only about half of the mass-loss debris will fall back onto the SMBH.
Since we can identify which SPH particles belong to the mass loss in the case of a head-on collision (Runs 1 & 3), here we will focus our analysis on the head-on collisions in both deep and gentle encounters. For grazing collisions, while the remnant properties differ, the mass-loss debris generally leaves the remnant’s Hill sphere quickly and is dominated by the SMBH’s gravitational influence. Therefore, when considering the fallback onto the SMBH, the results for grazing collisions are similar to those for head-on collisions.
Figure 10: Probability density function (PDF) of the orbital energy of the mass-loss debris relative to the SMBH after head-on collisions in Runs 1a and 3. The results are scaled by $\varepsilon_{\rm col}$, as defined in Eq. (13).
We begin by analyzing the orbital energy of the mass-loss debris relative to the SMBH, with the results shown in Figure 10. It is clear that the energy distribution of the debris produced by stellar collisions is quite different from that of the TDE debris. Here, we scale the orbital energy of the debris relative to the SMBH by the tidal energy of the SMBH at pericenter during a deep encounter:
$\varepsilon_{\rm col}\equiv\frac{GM_{\rm BH}R_{\star}}{r_{p}^{2}}=\frac{GM_{\rm BH}R_{\star}}{{r_{\rm tide}^{\star}}^{2}}\beta_{\star}^{2}.$ (13)
Although the collision in a deep encounter does not occur exactly at the pericenter, and this energy scale is less justified in the case of a gentle encounter, Figure 10 shows that Eq. (13) works well when describing the energy distribution, especially in terms of the peak value.
From the Figure 10, we see that the debris energy distribution exhibits a bimodal structure, with a relatively long tail. The bimodal feature arises because the orbital energy of the merger remnant relative to the SMBH is nearly zero, so most particles with nearly zero orbital energy remain bound to the merger remnant. The long tail indicates that many debris particles have relatively high launch velocities relative to the merger remnant. This is due to the shock waves fand fluid forces present during the collision, allowing some debris to achieve higher energies relative to the merger remnant. We also see that in our representative examples, the head-on collision in the deep encounter (Run 1a) results in a wider energy distribution compared to the head-on collision in the gentle encounter (Run 3).
For the mass-loss debris that falls back onto the SMBH, we are interested in their fallback rate, $\dot{M}(t)$, as it is strongly correlated with the observable light curve. Using the method for calculating $\dot{M}(t)$ described in Section 2.3, we compute the fallback time for each SPH particle identified as mass loss. The results are shown in Figure 11.
Figure 11: Fallback accretion rate, $\dot{M}(t)$, onto the SMBH from Runs 1 and 3. The different colored lines represent the results for different particle numbers ($N=5,1\times 10^{5}$ per star, Run 1a and 1b). To convert to physical units, here we set $M_{\rm BH}=10^{6}M_{\odot}$, $M_{\star}=1M_{\odot}$ and $R_{\star}=1R_{\odot}$. The black dashed line shows the standard mass fallback rate for a $1M_{\odot}$ star undergoing a TDE, while the red dashed line represents a power-law curve that matches the late-time fallback rate. The dotted line indicates the Eddington limit for a $10^{6}M_{\odot}$ black hole, assuming an accretion efficiency of $\epsilon=0.1$. Here, $t=0$ corresponds to the moment of stellar collision.
In this figure, we also compare the fallback rates for simulations with different particle numbers ($N=5,1\times 10^{5}$ per star) and contrast them with the standard mass fallback rate from the TDE of a $1M_{\odot}$ star, as well as the Eddington limit for a $10^{6}M_{\odot}$ SMBH, assuming an accretion efficiency of $\epsilon=0.1$.
The fallback rate curves of the mass-loss debris differ significantly from the standard TDE curve. These curves exhibit a slow rise over an extended period, remain relatively steady, and then follow a power-law decline similar to that of a TDE. The relationship between the curve shape and $\beta_{b}$ is not very clear. The peak fallback rate is much lower than that of a TDE, as the total fallback mass is only a few percent of the fallback mass in a TDE for a star of the same mass. However, the minimum fallback time for the debris is significantly shorter compared to the TDE case, particularly for deep encounters. In gentle encounters, this effect is less pronounced, as the collision occurs long after the pericenter passage. Despite comparable total mass loss, there is a significant difference in the magnitude of the peak fallback rates. In the deep encounter (Run 1), the collision occurs very close to the pericenter, causing the debris to fall back onto the SMBH more quickly. This results in a substantially higher peak rate, which remains above the Eddington limit for several hundred days. In contrast, for the gentle encounter (Run 3), the fallback occurs much more slowly, and the curve consistently stays below the Eddington limit.
After the mass-loss debris falls back to the pericenter around the SMBH, we expect it to undergo orbital circularization and disk formation, and eventually accrete onto the SMBH, analogous to that of TDE debris (Bonnerot & Stone, 2021).
However, since the “cloud” of debris produced by stellar collisions differs significantly from the TDE stream, the energy dissipation process may also differ substantially. A quantitative calculation of the light curve of BH accretion caused by stellar collisions is beyond the scope of this paper. Nevertheless, we can provide some qualitative insights. Due to the large width of the debris “cloud”, the compression effects during fallback onto the SMBH are expected to be more pronounced compared to the TDE streams, leading to greater energy dissipation from nozzle shocks. The inclination range of the debris produced by collisions is also broader than in TDEs. As a result, the density at pericenter may be lower than in the case of coplanar debris, but the intersection of orbits with different inclinations near pericenter could result in additional energy dissipation.
We can estimate the approximate size of the final disk by calculating the pericenter distances of the mass-loss SPH particles, based on their angular momenta (relative to the SMBH) after the collision. The results are shown in Figure 12.
Figure 12: Probability density function (PDF) of the pericenter distance ($r_{p}$) of the mass-loss debris relative to the SMBH in Runs 1a and 3. The results are scaled by the initial pericenter distance of the binary orbit relative to the SMBH $(r_{p0})$. The orange line represents all mass-loss debris, while the blue line represents the subset of debris that will fall back onto the SMBH.
We see that after the deep encounter head-on collision (Run 1a), the angular momentum distribution of the debris exhibits only about $5\%$ variation compared to the initial angular momentum, with the fallback debris primarily concentrated in the region where just inside $r_{p0}$ (the initial pericenter distance of the binary). In contrast, during the gentle encounter head-on collision, the angular momentum distribution of the debris varies by about $40\%$ compared to the initial value. The difference in these distributions arises because deep encounter collisions occur very close to the SMBH, where the orbital velocity is much greater than the velocity of the debris relative to the remnant. As a result, the debris angular momentum remains largely unaffected. On the other hand, gentle encounter collisions occur far from $r_{p0}$, at the location where the orbital velocity is already relatively small, making the angular momentum more susceptible to change.
4 Summary and Discussion
4.1 Summary of Key Results
In this paper, we have conducted a detailed study of the stellar collisions induced by the close encounter between a stellar binary and a SMBH. The binary, with a total mass $m_{12}$, an initial circular orbit, and a semi-major axis $a_{b}$, approaches the SMBH (mass $M_{\rm BH}$) on a parabolic orbit, with a pericenter distance $r_{p}$ from the SMBH. Using SPH simulations, we characterize key properties of stellar collisions and their potential observational consequences.
Our main findings focus on four representative cases involving both head-on and grazing collisions in gentle ($\beta_{b}\lesssim 1$) and deep ($\beta_{b}\gtrsim 1$) encounters (Table 1). We analyze the mass loss fraction, the density structure of the merger remnant(s), and the mass fallback rate.
  1. 1.
    Mass loss and remnant(s). Since the collision velocity is comparable to the stellar escape velocity (see Table 1 and Figure 9 in Paper I), some mass loss is expected during the collision. Moreover, depending on the impact parameter (e.g., head-on or grazing collisions), the outcome of the collision may vary. In the case of head-on collisions, the mass loss fraction is approximately $5\%$ (Figure 5), and such collisions consistently produce a single merger remnant. In contrast, grazing collisions exhibit more diverse outcomes. For gentle encounters, while the first collision may not immediately result in a single merger remnant, the binary remains gravitationally bound; the perturbed stars undergo multiple subsequent collisions (Figure 8), ultimately forming a single remnant with minimal mass loss ($\lesssim 1\%$). However, in deep encounters, where the binary system is already disrupted, if the first collision fails to produce a merger, two remnants will form, and the mass loss fraction is relatively higher (around $4\%$).
  2. 2.
    Density structure of merger remnant(s). The remnant star formed after a collision exhibits significant differences compared to the original main-sequence star (modeled here as a $\gamma=5/3$ polytrope). The merger remnant features a much higher central density and a highly extended envelope, with approximately $5\%$ of its mass residing beyond five times the original radius (Figs. 6, 9). This extended structure makes the merger remnant more susceptible to partial tidal disruption when it returns to the pericenter around the SMBH. If the two stars do not merge but instead undergo significant dynamical perturbations, the outcomes differ substantially (upper row of Figure 9). These perturbed stars not only develop an extended envelope but also experience a substantial reduction in the central density. Analogous to the Hills mechanism, one of the stars often remains bound to the SMBH. In this case, this perturbed star is also significantly more susceptible to tidal disruption.
  3. 3.
    Orbital property of mass-loss debris and mass fallback rate. Nearly half of the mass-loss debris produced during the collision is expected to fall onto the SMBH, undergoing a process analogous to a TDE. From the simulation snapshots (Figs. 1-4), we observe that the “debris cloud” generated by the collision exhibits a morphology distinctly different from the “debris stream” typical of TDEs. Furthermore, an analysis of the binding energy distribution of the debris (Figure 10) reveals significant deviations from the top-hat distribution commonly observed in stellar TDEs. Since the collision does not necessarily occur at the pericenter, unlike in the standard TDE, the resulting mass fallback accretion rate also differs (Figure 11). Although the total mass loss represents only a small fraction of the stellar mass and thus results in a relatively low magnitude of fallback accretion, the shape of the fallback rate curve is notably distinct from that of TDEs. Specifically, it remains approximately constant for an extended period ($\gtrsim 100$ days) before starting to decline (for $M_{\rm BH}=10^{6}M_{\odot}$ and solar-type stars). Additionally, the angular momentum distribution of the mass-loss debris (represented by the debris’ pericenter distance $r_{p}$, see Figure 12) suggests that the resulting accretion disk extend to the original pericenter distance.
4.2 Parameter Space and Comments on the Results of Paper I
Figure 13: Probability density function (PDF) of the relative orbital angular momentum (left panel) and impact parameter $b_{12}/r_{12}$ (right panel) when two stars contact each other. The blue and orange solid lines represent the results for two different values of $\beta_{b}$: gentle encounter ($\beta_{b}=0.6$) and deep encounter ($\beta_{b}=5$). The red dashed line in the left panel represents the typical relative angular momentum for grazing collisions, where the contact velocity is the escape velocity $v_{\rm esc}=\sqrt{2Gm_{12}/R_{\rm col}}$ and impact parameter is $b_{12}=R_{\rm col}$.
In this paper, we have carried out hydrodynamical simulations in four stellar collision cases (see Table 1), involving both gentle ($\beta_{b}=0.6$) and deep ($\beta_{b}=5$) encounters, leading to nearly head-on ($b_{12}/r_{12}\lesssim 0.06$) and more grazing ($b_{12}/r_{12}\simeq 0.5-0.9$) collisions. How representative are these cases? In Paper I, we considered a wide range of parameter space using N-body integrations, and calculated the probabilities of stellar collisions in binary-SMBH encounters; we adopted the inelastic collision assumption, which implies that all collisions lead to merger with no mass loss. Obviously, the collision probabilities and the impact velocity distribution obtained in Paper I remain valid. The outcome of a collision, on the other hand, depends strongly on the collisional impact parameter ($b_{12}/r_{12}$).
To determine the range of outcomes of stellar collisions in binary-SMBH encounters, we show in Figure 13 the probability distributions of the relative orbital angular momentum $L$ (left panel) and impact parameter $b_{12}/r_{12}$ (right panel) of the two colliding stars for both gentle and deep encounters, based on the simulation data of Paper I. We see that most collisions are grazing collisions, characterized by relatively large $b_{12}/r_{12}$. The hydrodynamical simulations reported in this paper show that for such grazing collisions, the outcomes differ between the encounter types: for gentle encounters, multiple collisions eventually lead to a single merger remnant; for deep encounters, the most common outcome is not a single merger remnant but rather two stars that have undergone significant dynamical perturbations.
Nevertheless, the two key outcomes that are potentially observable – mass loss and fallback accretion, and the formation of puffier stars – remain largely unaffected. Even in cases with two perturbed stars, a fraction of the mass loss can still be accreted onto the SMBH, producing a flare, and the stars involved are significantly puffier than their initial states. Moreover, in such cases, one of the perturbed stars remains bound to the SMBH, typically on a shorter orbital period compared to the single merger remnant (see the comparison between Sections 6.2.1 and 6.2.2 in Paper I). This further supports the idea that stellar collisions induced by binary-SMBH encounters can enhance the likelihood of repeated partial tidal disruption events.
This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.